Forty decreased by the number m is 24.5?
The sumof one third of the number s and 12 is 32?
Four less than twice a number is nine?
Three times two dcreased by x is eight?
Three diminished by twice a number is eight?



Answer by rfer(16322)

40-m=24.5
1/3s+12=32
2x-4=9
3*2-x=8
3-2x=9